# Uncertainty Calibration of Multi-Label Bird Sound Classifiers

### v1

@@ 3 EXPERIMENTS
Table 2: Overview of key characteristics of bird sound classification models. For each model, we indicate the year of release, the number of classes the model is trained to classify, training method—supervised learning (SL) or self‐supervised learning (SSL)—as well as the architecture, number of parameters, embedding size, input duration (in seconds), input type and sample rate (in Hz).
@@ 3.1.1 Datasets
While many datasets in bioacoustics exist, BirdSet [Rauch et al., 2024] is not only widely used in the field [van Merriënboer et al., 2025, Miron et al., 2025, Heinrich et al., 2025, Schwinger et al., 2025], but particularly well-suited for our experiments due to its multi-label nature and practical relevance. BirdSet consists of eight test datasets sourced from PAMdeployments, covering a range of habitats and recording conditions. These test datasets are fully annotated by experts, while the training data is sourced from the public Xeno-Canto database. Table 1 gives an overview of the datasets.
@@ 3.1.2 Models – Bird Sound Classifiers
ConvNeXtBS is a ConvNeXt [Liu et al., 2022] encoder with a linear classification head. The model weights were initialised using an ImageNet pretrained checkpoint and further trained via supervised learning on the BirdSet XCL training data. Extensive augmentations are used to allow the model to generalise better for the task of bird sound classification in PAMscenarios, see [Rauch et al., 2024] for more details.
@@
AudioProtoPNet [Heinrich et al., 2025] is an interpretable deep learning model for multi-label bird sound classification, adapted from the Prototypical Part Network (ProtoPNet) architecture [Chen et al., 2019]. It leverages a ConvNeXt backbone to extract embeddings from input spectrograms and learns prototypical patterns for each bird species in the embedding space. During inference, the model classifies recordings by comparing them to learned prototypes, providing both accurate predictions and interpretable explanations for its decisions. Like ConvNeXtBS, it is trained via supervision on the BirdSet XCL dataset. We use the largest version with 20 prototypes per class.
@@
BirdMAE [Rauch et al., 2025] is a domain-specialised masked autoencoder (MAE) designed for bird sound classification. Unlike general-purpose audio MAEs pretrained on AudioSet, BirdMAE is pretrained on the BirdSet XCL dataset to better capture the fine-grained acoustic characteristics of avian bioacoustics. The model supports both fine-tuning and parameter-efficient prototypical probing, where frozen representations are leveraged for downstream tasks. For this paper, we fine-tuned the base checkpoint on the BirdSet XCL training data using a linear classification head.
@@
Perch v2 [van Merriënboer et al., 2025] is a state-of-the-art model for bioacoustic sound classification. It builds upon the original Perch architecture (EfficientNetB3) by incorporating additional training data and advanced augmentation techniques, resulting in improved performance across various evaluation metrics.
We excluded BirdNET [Kahl et al., 2021] from our experiments as it used the BirdSet datasets for training and validation, which would bias our calibration assessment.
@@ 3.1.3 Metrics
cmAP is obtained by computing the average precision (AP) for each class $c$ independently and then averaging these values across all $C$ classes, yielding a macro-average [Rauch et al., 2024]. It can be computed by:
@@
This metric evaluates the model’s ability to consistently rank positive instances above negative ones across all possible decision thresholds, providing a threshold-independent measure of overall performance. Nonetheless, it can become noisy for classes with few positive labels, which reduces its comparability across datasets [Denton et al., 2022].
In a multi-label setting, reliability diagrams can be computed for each class for each class individually and aggregated for simpler visualisation.
@@
ECE quantifies the difference between accuracy and confidences. The predictions are partitioned into $M$ bins, the error can be computed as follows:
@@
As the ECE only quantifies the calibration error and gives no information on whether the model is over or underconfident, the MCS [Ao et al., 2023] computes the difference between confidence and accuracy in each bin. A perfectly calibrated model has an MCS of 0 while under-confident models have a negative score and overconfident models a positive score. The ECE equation is adapted to:
@@
As the MCS computes a average of the difference between $\text{conf}(B_{m})$ and $\text{acc}(B_{m})$ over and underconfident bins can cancel each other out. Consequently, the MCS omits information if the model is miscalibrated in multiple regions. Therefore, we propose to additionally evaluate:
@@
These metrics quantify the extent of over and underconfidence separately, providing a more nuanced view of calibration performance.
@@ 3.2 Benchmarking the Calibration of Bird Sound Classifiers
In this section, we benchmark the calibration performance of multiple models using eight strongly labelled BirdSet evaluation datasets. We examine calibration at different levels of granularity, from global and per-dataset assessments to per-class and class-frequency, to determine how calibration varies across domains and classes.
@@ 3.2.1 Experiment: Global and Per-dataset calibration
All test samples are collected to report the global performance for each model, neglecting the origin of the samples. Additionally domain-specific calibration performance is investigated across individual BirdSet evaluation datasets. This experiment investigates calibration variability between domains (datasets) and how that compares to calibration metrics when aggregating across all domains.
@@ 3.2.2 Experiment: Per-class calibration
We evaluate the per-class ECE given the number of samples per class, investigating calibration characteristics for each model on a class basis. This experiment investigates the generalisability of model calibration metrics across imbalanced classes by evaluating how calibration metrics change when a class has more samples.
@@ 3.3 Post hoc Methods to Improve Calibration
We also investigate whether calibration can be improved using the compute-efficient post hoc methods TSand PS. Both methods need specific data to optimise the scaling parameters. We experiment in two settings: (i) utilise the BirdSet POW dataset as a calibration set and (ii) leverage the first 10 minutes of each dataset.
As the classes of the POW dataset differ from the test datasets, only global scaling parameters (temperature or slope and bias) can be fitted, whereas using a part of the test dataset enables fitting per-class parameters. The test dataset is slightly reduced in this case, making a direct comparison between the two approaches difficult. Still, in practice it is often possible to label a small part of the test dataset. In both settings and for each method, we optimise the parameters using Adam [Kingma, 2014] with a learning rate of $0.001$ for 1000 steps.
@@ 4 Results and Discussion
Table 3: The models AudioProtoPNet, BirdMAE, ConvNeXtBS and Perch v2 are evaluated on the 8 BirdSet test sets with the metrics cmAP, ECE, MCS, OCS and UCS. The ”All” column reflects the class-wise average independent of dataset. Strongest performance is highlighted in bold. Arrows next to the metrics indicate what qualifies as good performance.
@@ 4.1.1 Global and per-dataset calibration
Table 3 summarises the results of the described metrics cmAP, ECE, MCS, OCS and UCS for each of the 8 BirdSet test sets as well as the class-wise average performance irrespective of dataset origin. All metrics are weighted by class frequency.
@@
Across all test samples, AudioProtoPNet achieves the strongest discriminative performance by cmAP with 41.36, followed by Perch v2 with 39.16. BirdMAE and ConvNeXtBS follow with 35.57 and 32.20, respectively. Both top-performing models set the highest scores on a per-dataset basis as well. Our results are in slight contradiction to those reported by Merriënboer et al. [van Merriënboer et al., 2025], since we use the larger version of AudioProtoPNet-20 and our BirdMAE model is fine-tuned in XCL rather than individually for each test dataset. Our BirdMAE checkpoint further utilises a simple classification head on top of the mean-pooled embeddings compared to the patch embeddings, which degrades performance.
@@
Considering the class-wise average independent of the test datasets in Table 3, AudioProtoPNet, which reaches the strongest performance by cmAP, is less calibrated than Perch v2 or ConvNeXtBS. Perch v2 appears to be calibrated best overall with an MCS of -0.91, suggesting slight underconfidence. ConvNeXtBS performs similar with an MCS of -1.09, also indicating slight underconfidence. In contrary, AudioProtoPNet and BirdMAE are overconfident with an MCS of 4.43 and 7.62, respectively. The misalignment of best performance by cmAP and MCS reiterates the importance of uncertainty calibration.
@@
Figure 3: Demonstrating model (ConvNeXt) calibration performance across individual BirdSet evaluation datasets and the combined performance when aggregated across all datasets.
@@
Considering ECE scores, Perch v2 and ConvNeXtBS achieve similar calibration scores with Perch v2 slightly outperforming ConvNeXtBS for all sets except for the NES test set. For all test sets except for POW and UHH, BirdMAE achieves the highest ECE score indicating poor calibration. For all models, the combined ECE, aggregated across all BirdSet evaluation datasets, is better than the average of all datasets. Per-dataset ECE values range from 1.37 for Perch v2 at SSW to 22.46 at POW. This demonstrates the importance of per-dataset evaluation, as calibration performance is potentially inflated when combining calibration metrics across distinct subdomains, masking the true calibration characteristics.
@@
When analysing the miscalibration using MCS, Perch v2 achieves the best scores for NBP, SSW and SNE, while now AudioProtoPNet outperforms Perch v2 on PER. Perch v2 and ConvNeXtBS are underconfident in all datasets whereas AudioProtoPNet and BirdMAE are overconfident except for POW and UHH or POW, respectively.
@@
The consistent behaviour of underconfidence is in line with the reliability diagram in Figure 1, in which Perch v2 and ConvNeXtBS exhibit values above the perfect calibration trajectory in all instances. OCS scores further confirm that Perch v2 and ConvNeXtBS achieve nearly ideal performance in respect to overconfidence. Subsequently, UCS scores carry the weight of the miscalibration for the two models. While Figure 1 shows performance as a function of predicted probabilities and therefore does not have to align with the MCS scores, it does indicate whether the model exhibits a general over or underconfident behaviour. AudioProtoPNet and BirdMAE achieve consistently positive MCS scores. Figure 1 shows that both models are overconfident for low predicted probabilities and underconfident for high predicted probabilities.
@@
Figure 3 shows a reliability diagram for ConvNeXtBS over all eight test sets and the combined set. Significant calibration variability across the evaluation datasets can be observed with most of the test set performances lying above the perfect calibration line, indicating underconfidence. Figure 3 also highlights the potential misalignment of MCS values and reliability diagrams. The reason for this is the very unequal distribution of samples in the predicted probability bins. The UHH test set produces a reliability diagram which seems to be very similar to that of the SSW test set; however, the MCS values reach -16.07 and -1.86, respectively, suggesting a large difference in calibration.
@@
Judging from the reliability diagram in Figure 3, ConvNeXt is overconfident for the datasets PER and NES which correspond to tropical locations, while the model is underconfident for other datasets. While beyond the scope of this paper, a more detailed investigation into the number of samples per bin for specific datasets could disentangle why the reliability diagram suggests overconfidence and the MCS values suggest underconfidence for the PER and NES datasets. In summary, a fine-grained evaluation of calibration metrics is a valuable tool to help understand model performance in different environments.
@@ 4.1.2 Per-class calibration
Figure 4 shows the ECE scores per class over the number of samples in each class. Different models are shown by colour. The well-performing models by ECE according to Table 3 are Perch v2 and ConvNeXtBS, which show very low ECE scores exhibiting a linear increase with the number of samples per class. AudioProtoPNet and BirdMAE, on the other hand, show less consistent ECE values and both models seem to exhibit a distinct split between many classes reaching low ECE values ($ECE<0.04$) and higher ECE values ($ECE>0.06$). For the largest class AudioProtoPNet and BirdMAE reach ECE values below Perch v2 and ConvNeXtBS.
The very large spread of ECE values for the underrepresented classes in AudioProtoPNet and BirdMAE provides an explanation for the poor performance per dataset in the previous section. Well-calibrated classes can be masked by poorly classified samples and result in low overall calibration performances. The distinct difference in class-wise performance could be influenced by some classes occurring only in very challenging noise conditions, resulting in an overall underperformance for that class. It further highlights that calibration scores can vary greatly between classes of very similar sizes.
@@
By contrast, AudioProtoPNet and BirdMAE display larger calibration variability between common and rare classes. For rare species, the reliability curves bend further away from the identity line at low probabilities (indicating stronger underconfidence), while at higher probabilities they get closer to the line (less underconfidence). This pattern yields positive MCS for the rare subset and increases OCS—consistent with their broader per-class ECE spread in Figure 4. In practice, this suggests that rare classes are particularly susceptible to overconfident predictions in these two models, which can be problematic for downstream thresholding and decision support.
@@
Table 4: Post hoc calibration results using temperature scaling (TS) and Platt scaling (PS) on BirdSet test sets measured with MCS. ’TS POW’ and ’PS POW’ indicate global scaling parameters fitted on the POW dataset, while ’TS per class’ and ’PS per class’ indicate per-class scaling parameters fitted on the first 10 minutes of each test dataset. The ’Base’ row reflect performance without post hoc calibration on the remaining test set. As a subscript the relative performance improvement is marked; for the Base row the absolute difference to the full test set is given. Best results are highlighted in bold.

### v2

@@ 3 EXPERIMENTS
Table 2: Overview of key characteristics of bird sound classification models. For each model, we indicate the year of release, the number of classes the model is trained to classify, training method—supervised learning (SL) or self‐supervised learning (SSL)—as well as the architecture, number of parameters, embedding size, input duration (in seconds), input type and sample rate (in kHz).
@@ 3.1.1 Datasets
While many datasets in bioacoustics exist, BirdSet (Rauch et al., 2025b, ) is not only widely used in the field (van Merriënboer et al.,, 2025; Miron et al.,, 2025; Heinrich et al.,, 2025; Schwinger et al.,, 2025), but also particularly well-suited for our experiments due to its multi-label nature and practical relevance. BirdSet consists of seven test datasets sourced from PAMdeployments, covering a range of habitats and recording conditions. These seven test datasets are fully annotated by experts, while the training data is sourced from the public Xeno-Canto database. Table 1 gives an overview of the datasets.
@@ 3.1.2 Models – Bird Sound Classifiers
ConvNeXtBS is a ConvNeXt (Liu et al.,, 2022) encoder with a linear classification head. The model weights were initialised using an ImageNet pretrained checkpoint and further trained via supervised learning on the BirdSet XCL training data. Extensive augmentations are used to allow the model to generalise better for the task of bird sound classification in PAMscenarios, see (Rauch et al., 2025b, ) for more details.
@@
AudioProtoPNet (Heinrich et al.,, 2025) is an interpretable deep learning model for multi-label bird sound classification, adapted from the Prototypical Part Network (ProtoPNet) architecture (Chen et al.,, 2019). It leverages a ConvNeXt backbone to extract embeddings from input spectrograms and learns prototypical patterns for each bird species in the embedding space. During inference, the model classifies recordings by comparing them to learned prototypes, providing both accurate predictions and interpretable explanations for its decisions. Like ConvNeXtBS, it is trained via supervision on the BirdSet XCL dataset. We use the largest version with 20 prototypes per class.
@@
BirdMAE (Rauch et al., 2025a, ) is a domain-specialised masked autoencoder (MAE) designed for bird sound classification. Unlike general-purpose audio MAEs pretrained on AudioSet, BirdMAE is pretrained on the BirdSet XCL dataset to better capture the fine-grained acoustic characteristics of avian bioacoustics. The model supports both fine-tuning and parameter-efficient prototypical probing, where frozen representations are leveraged for downstream tasks. For this paper, we fine-tuned the base checkpoint on the BirdSet XCL training data using a linear classification head.
@@
Perch v2 (van Merriënboer et al.,, 2025) is a state-of-the-art model for bioacoustic sound classification. It builds upon the original Perch architecture (EfficientNet-B3) by incorporating additional training data and advanced augmentation techniques, resulting in improved performance across various evaluation metrics.
Table 2 provides an overview of key model properties. We excluded BirdNET (Kahl et al.,, 2021) from our experiments as its training data overlaps with the BirdSet evaluation datasets, which would bias our calibration assessment.
@@ 3.1.3 Metrics
cmAP is obtained by computing the average precision (AP) for each class $c$ independently and then averaging these values across all $C$ classes, yielding a macro-average (Rauch et al., 2025b, ). It can be computed by:
@@
This metric evaluates the model’s ability to consistently rank positive instances above negative ones across all possible decision thresholds, providing a threshold-independent measure of overall performance. Nonetheless, it can become noisy for classes with few positive labels, which reduces its comparability across datasets (Denton et al.,, 2022).
In a multi-label setting, reliability diagrams can be computed for each class individually and aggregated for simpler visualisation.
@@
ECE quantifies the difference between accuracy and confidences. The predictions are partitioned into $M$ bins (we use $M=15$ in our experiments). The error can be computed as follows:
@@
As ECE only quantifies the calibration error and gives no information on whether the model is over- or underconfident, the MCS (Ao et al.,, 2023) computes the difference between confidence and accuracy in each bin. A perfectly calibrated model has an MCS of 0 while under-confident models have a negative score and overconfident models a positive score. The ECE equation is adapted to:
@@
As the MCS computes an average of the difference between $\text{conf}(B_{m})$ and $\text{acc}(B_{m})$, over- and underconfident bins can cancel each other out. Consequently, the MCS loses information if the model is miscalibrated in multiple regions. Therefore, we additionally evaluate:
@@
These metrics quantify the extent of over- and underconfidence separately, providing a more nuanced view of calibration performance.
@@ 3.2 Benchmarking the Calibration of Bird Sound Classifiers
In this section, we benchmark the calibration performance of multiple models using seven strongly labelled BirdSet test datasets and the POW validation set. We examine calibration at different levels of granularity, from global and per-dataset assessments to per-class and class-frequency levels, to determine how calibration varies across domains and classes.
@@ 3.2.1 Experiment: Global and Per-dataset calibration
All test samples are collected to report the global performance for each model, neglecting the origin of the samples. Additionally, domain-specific calibration performance is investigated across individual BirdSet evaluation datasets. This experiment investigates calibration variability between domains (datasets) and how that compares to calibration metrics when aggregating across all domains.
@@ 3.2.2 Experiment: Per-class calibration
We evaluate the per-class ECE given the number of samples per class, investigating calibration characteristics for each model on a class basis. This experiment investigates the generalisability of model calibration across imbalanced classes by evaluating how calibration metrics change when a class has more samples.
@@ 3.3 Post hoc Methods to Improve Calibration
We also investigate whether calibration can be improved using the compute-efficient post hoc methods TS (Guo et al.,, 2017) and PS (Platt,, 1999). Both methods need specific data to optimise the scaling parameters. We experiment in two settings: (i) utilise the BirdSet POW dataset as a calibration set and (ii) leverage the first 10 minutes of each dataset.
As the classes of the POW dataset differ from the test datasets, only global scaling parameters (temperature, or temperature and bias) can be fitted, whereas using a part of the test dataset enables fitting per-class parameters. The test dataset is slightly reduced in this case, making a direct comparison between the two approaches difficult. Still, in practice it is often possible to label a small part of the test dataset. In both settings and for each method, we optimise the parameters using Adam (Kingma and Ba,, 2015) with a learning rate of $0.001$ for 1000 steps.
@@ 4 RESULTS AND DISCUSSION
Table 3: The models AudioProtoPNet, BirdMAE, ConvNeXtBS and Perch v2 are evaluated on the seven BirdSet test sets and the POW validation set with the metrics cmAP, ECE, MCS, OCS and UCS. The ”All” column reflects the class-wise average independent of dataset. Strongest performance is highlighted in bold. Arrows next to the metrics indicate what qualifies as good performance.
@@ 4.1.1 Global and per-dataset calibration
Table 3 summarises the results of the described metrics cmAP, ECE, MCS, OCS and UCS for each of the seven BirdSet test sets and the POW validation set as well as the class-wise average performance irrespective of dataset origin. All metrics are weighted by class frequency.
@@
Across all test samples, AudioProtoPNet achieves the strongest discriminative performance by cmAP with 41.36, followed by Perch v2 with 39.16. BirdMAE and ConvNeXtBS follow with 35.57 and 32.20, respectively. Both top-performing models set the highest scores on a per-dataset basis as well. Our results slightly differ from those reported by van Merriënboer et al., (2025), since we use the larger version of AudioProtoPNet-20 and our BirdMAE model is fine-tuned on XCL rather than individually for each test dataset. Our BirdMAE checkpoint further utilises a simple classification head on top of the mean-pooled embeddings rather than the patch embeddings, which degrades performance.
@@
Considering the class-wise average independent of the test datasets in Table 3, AudioProtoPNet, which reaches the strongest performance by cmAP, is less calibrated than Perch v2 or ConvNeXtBS. Perch v2 appears to be calibrated best overall with an MCS of -0.91, suggesting slight underconfidence. ConvNeXtBS performs similarly with an MCS of -1.09, also indicating slight underconfidence. In contrast, AudioProtoPNet and BirdMAE are overconfident with an MCS of 4.43 and 7.01, respectively. The misalignment of best performance by cmAP and MCS reiterates the importance of uncertainty calibration.
@@
Figure 3: Demonstrating model (ConvNeXtBS) calibration performance across individual BirdSet evaluation datasets and the combined performance when aggregated across all datasets.
@@
Considering ECE scores, Perch v2 and ConvNeXtBS achieve similar calibration scores with Perch v2 slightly outperforming ConvNeXtBS for all sets except for the NES test set. For all test sets except for POW and UHH, BirdMAE achieves the highest ECE score, indicating poor calibration. For all models, the combined ECE, aggregated across all BirdSet evaluation datasets, is better than the average of all datasets. Per-dataset ECE values for Perch v2 range from 1.37 at SSW to 22.46 at POW. This demonstrates the importance of per-dataset evaluation, as calibration performance is potentially inflated when combining calibration metrics across distinct subdomains, masking the true calibration characteristics.
@@
When analysing the miscalibration using MCS, Perch v2 achieves the best scores for NBP, SSW and SNE, while now AudioProtoPNet outperforms Perch v2 on PER. Perch v2 and ConvNeXtBS are underconfident in all datasets whereas AudioProtoPNet and BirdMAE are overconfident except for POW and UHH (AudioProtoPNet) or POW (BirdMAE), respectively.
@@
The consistent behaviour of underconfidence is in line with the reliability diagram in Figure 1, in which Perch v2 and ConvNeXtBS exhibit values above the perfect calibration trajectory in all instances. OCS scores further confirm that Perch v2 and ConvNeXtBS achieve nearly ideal performance with respect to overconfidence. Subsequently, UCS scores carry the weight of the miscalibration for the two models. While Figure 1 shows performance as a function of predicted probabilities and therefore does not have to align with the MCS scores, it does indicate whether the model exhibits a general over- or underconfident behaviour. AudioProtoPNet and BirdMAE achieve consistently positive MCS scores. Figure 1 shows that both models are overconfident for low predicted probabilities and underconfident for high predicted probabilities.
@@
Figure 3 shows a reliability diagram for ConvNeXtBS over the seven test sets, the POW validation set, and the combined set. Significant calibration variability across the evaluation datasets can be observed with most of the test set performances lying above the perfect calibration line, indicating underconfidence. Figure 3 also highlights the potential misalignment of MCS values and reliability diagrams. The reason for this is the very unequal distribution of samples in the predicted probability bins. The UHH test set produces a reliability diagram which seems to be very similar to that of the SSW test set; however, the MCS values reach -16.07 and -1.86, respectively, suggesting a large difference in calibration.
@@
Judging from the reliability diagram in Figure 3, ConvNeXtBS is overconfident for the datasets PER and NES, which correspond to tropical locations, while the model is underconfident for other datasets. While beyond the scope of this paper, a more detailed investigation into the number of samples per bin for specific datasets could disentangle why the reliability diagram suggests overconfidence and the MCS values suggest underconfidence for the PER and NES datasets. In summary, a fine-grained evaluation of calibration metrics is a valuable tool to help understand model performance in different environments.
@@ 4.1.2 Per-class calibration
Figure 4 shows the ECE scores per class over the number of samples in each class. Different models are shown by colour. The well-performing models by ECE according to Table 3 are Perch v2 and ConvNeXtBS, which show very low ECE scores that exhibit a linear increase with the number of samples per class. AudioProtoPNet and BirdMAE, on the other hand, show less consistent ECE values and both models seem to exhibit a distinct split between many classes reaching low ECE values ($ECE<0.04$) and higher ECE values ($ECE>0.06$). For the largest class, AudioProtoPNet and BirdMAE reach ECE values below Perch v2 and ConvNeXtBS.
The very large spread of ECE values for the underrepresented classes in AudioProtoPNet and BirdMAE provides an explanation for the poor performance per dataset in the previous section. Well-calibrated classes can be masked by poorly calibrated classes, resulting in low overall calibration performance. The distinct difference in class-wise performance could be influenced by some classes occurring only in very challenging noise conditions, resulting in an overall underperformance for that class. It further highlights that calibration scores can vary greatly between classes of very similar sizes.
@@
By contrast, AudioProtoPNet and BirdMAE display larger calibration variability between common and rare classes. For rare species, the reliability curves bend further away from the identity line at low probabilities (indicating stronger overconfidence), while at higher probabilities they get closer to the line (less underconfidence). This pattern yields positive MCS for the rare subset and increases OCS—consistent with their broader per-class ECE spread in Figure 4. In practice, this suggests that rare classes are particularly susceptible to overconfident predictions in these two models, which can be problematic for downstream thresholding and decision support.
We note that we define rare species solely from the test sets; further investigation could take the train set distribution into consideration. In addition, the used metrics are sensitive to individual samples, when only a few instances per class are available.
@@
Table 4: Post hoc calibration results using temperature scaling (TS) and Platt scaling (PS) on BirdSet test sets measured with MCS. ’TS POW’ and ’PS POW’ indicate global scaling parameters fitted on the POW dataset, while ’TS per class’ and ’PS per class’ indicate per-class scaling parameters fitted on the first 10 minutes of each test dataset. The ’Base’ row reflects performance without post hoc calibration on the remaining test set. As a subscript the relative performance improvement is marked; for the ’Base’ row the absolute difference to the full test set is given. Best results are highlighted in bold.